OpenStudy (anonymous):

Determine whether the improper integral converges or diverges.∫[0 on top &-∞on the bottom] e^x dx

OpenStudy (nikita2):

cobverges

OpenStudy (nikita2):

this int e^x = 1

myininaya (myininaya):

\[\int\limits_{-\infty}^{0}e^x dx=\lim_{b \rightarrow -\infty}\int\limits_{b}^{0}e^x dx=\lim_{b \rightarrow -\infty} e^x, x=b..0\] \[\lim_{b \rightarrow -\infty}(e^0-e^b)=1-\lim_{b \rightarrow -\infty}e^b=1-0=1\]

myininaya (myininaya):

gj nikita

myininaya (myininaya):

i hope you know what i mean when i e^x, x=b..0 that means i'm just fixing to evaluate e^0-e^b okay?

OpenStudy (anonymous):

ok.understood!

myininaya (myininaya):

i did that last time on a previous problem with you tonight

OpenStudy (anonymous):

yes. thanks. which country r u from?

myininaya (myininaya):

i can't use the other symbol or if i can i don't know how so i made up my own notation for it

myininaya (myininaya):

usa

OpenStudy (anonymous):

ok. i'm from Singapore.

myininaya (myininaya):

i heard of your country but know little about it lol i'm just a dumb american

OpenStudy (anonymous):

haha

OpenStudy (nikita2):

Oh guys! I'm from Russia)

OpenStudy (anonymous):

oh i see

myininaya (myininaya):

nikita is that your real pic?

myininaya (myininaya):

pic=picture

myininaya (myininaya):

you knew that lol

OpenStudy (nikita2):

hm? yes)

myininaya (myininaya):

you look cute

myininaya (myininaya):

and so do you joe lol

OpenStudy (anonymous):

lol thanks :)

OpenStudy (anonymous):

i need to buy a calculus book or something, im very very very very weak in calculus :(